### v7

@@ 1 Introduction
The next section presents a model of strategic network formation and defines network moments. In § 3, we state high-level conditions for a CLT and outline its proof. Readers interested in the low-level CLT for strategic network formation may skip § 3 for § 4, which presents primitive sufficient conditions. We also outline a general methodology for verifying the high-level condition that can be applied to other network models. In § 5, we discuss practical inference procedures. We present results from a simulation study in § 6, and § 7 concludes. All proofs are stated in the supplemental appendix.
@@ 2.1 Homophily and Sparsity
For any $\delta\in\mathbb{R}_{+}$, $\bar{V}(\delta,\cdot)$ is invertible in its second argument, and its inverse $\bar{V}^{-1}(\delta,\cdot)$ satisfies $\limsup_{\delta\rightarrow\infty}\delta^{-1}\log\bar{\Phi}_{\zeta}(\bar{V}^{-1}(\delta,0))<0$. Furthermore, there exists $\kappa>0$ such that, for any $n\in\mathbb{N}$,
@@ 2.2 Strategic Interactions and Equilibrium Selection
where $\{S_{n}\}_{n\in\mathbb{N}}$ is a sequence of permutation-invariant functions. [^1] For any injection $\pi\colon\mathcal{N}_{n}\rightarrow\mathcal{N}_{n}$, we abuse notation by defining $\pi(r_{n}^{-1}\bm{T}_{n})=((r_{n}^{-1}X_{\pi(i)},Z_{\pi(i)}))_{i\in\mathcal{N}_{n}}$ and $\pi(\bm{A})=(A_{\pi(i)\pi(j)})_{i,j}$. We say $S_{n}(\cdot)$ is permutation-invariant if $S_{n}(i,j,r_{n}^{-1}\bm{T}_{n},\bm{A})=S_{n}(\pi(i),\pi(j),\pi(r_{n}^{-1}\bm{T}_{n}),\pi(\bm{A}))$. This means that the function does not directly depend on node labels which have no economic content in our model.
@@ 3.2 Outline of Proof
We prove (13) by adapting a theorem due to Penrose and Yukich (2008) (see our Theorem SA.1.1). We proceed by first partitioning the support of $X_{1}$ into cubes $Q_{1},\dots,Q_{V_{n}}$ of slowly growing volume and representing the moment $n^{-1}\sum_{i=1}^{n}\psi_{i}(\mathcal{N}_{N_{n}})$ as a double sum over cubes and nodes within cubes $n^{-1}\sum_{i=1}^{V_{n}}\sum_{j=1}^{N_{i}}\psi_{ij}(\mathcal{N}_{N_{n}})$, where $N_{i}$ is the number of nodes positioned in $Q_{i}$ and $ij$ indexes the $j$th node in $Q_{i}$. Spatial independence of the Poisson process implies independence of node positions across cubes. However, node statistics are complex functionals of the process and hence are not generally independent across cubes.
Since stabilization implies that node statistics $\psi_{ij}(\mathcal{N}_{n})$ are primarily determined by nodes relatively proximate to $ij$, and hence cubes relatively proximate $Q_{i}$, it can be shown that the dependence structure of $\{\sum_{j=1}^{N_{i}}\psi_{ij}(\mathcal{N}_{N_{n}})\}_{i=1}^{V_{n}}$ is “approximately” characterized by a certain “dependency graph.” This is a network in which, roughly speaking, two observations are linked if and only if they are dependent. In our case, observations are cubes, which we connect if and only if they are relatively proximate. Careful construction of the cubes ensures a small approximation error for the dependency graph characterization, and applying a CLT for dependency graphs (Chen and Shao, 2004) delivers the result. For precise details, see § SA.1.1.
@@ 4.1 Strategic Neighborhood
$D_{ij}=\bm{1}\big{\{}-\theta_{2}<\theta_{1}-\theta_{3}(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert)^{2}+\zeta_{12}\leq 0\big{\}}.$
@@ 4.2.1 Best-Response Chains
The previous example pertains to interventions that perturb non-robust potential links, specifically $A_{56}$. Interventions that perturb robust links, such as $A_{67}$, trigger best-response dynamics in all strategic neighborhoods containing the link, which would be the two right-most components in Figure 1.
@@ 4.3 Strength of Interactions
$h^{*}(x,z)$ $=\kappa\int_{\mathbb{R}^{2}}\left[\Phi(\theta_{1}+\theta_{2}-\theta_{3}\lVert x\rVert^{2})-\Phi(\theta_{1}-\theta_{3}\lVert x\rVert^{2})\right]\,\text{d}x$
@@
$h^{*}(x^{\prime},z^{\prime})=\frac{\kappa\pi}{\theta_{3}}\big{(}q(\theta_{1}+\theta_{2})-q(\theta_{1})\big{)}.$
@@ 4.4 Coordination
However, if $H$ is specifically a strategic neighborhood, then under Assumption 2, its subnetwork is in fact pairwise stable regardless of the state of the ambient network. Consider Figure 1 and a counterfactual intervention that removes all nodes from the network outside of $C_{5}^{+}=\{2,\ldots,7\}$. The intervention has no impact on the pairwise stability of $A_{23}$ and $A_{67}$ since these are robust links. It also has no impact on the pairwise stability of non-robust potential links between nodes in $C_{5}^{+}$, for example $A_{45}$ since neither nodes 4 or 5 are linked to $\mathcal{N}_{n}\backslash C_{5}^{+}$ in $\bm{A}$ and strategic interactions are local by Assumption 2. That is, the non-robust potential links in $C_{5}^{+}$ are separated from $\mathcal{N}_{n}\backslash C_{5}^{+}$ by a “buffer” of robust links, so the intervention does not trigger a chain of best-responses that affects $\bm{A}_{C_{5}^{+}}$. This illustrates the following result.
@@
Recall that components of $\bm{D}$ do partition $\mathcal{N}_{n}$, while strategic neighborhoods are obtained by adding nodes that are robustly linked to components. Then the link between any node pair that lies in multiple strategic neighborhoods must necessarily be robust and therefore have the same realization under any pairwise stable equilibrium. That is, any pairwise stable subnetwork on the left strategic neighborhood in Figure 1 sets $A_{23}=1$, as does any pairwise stable subnetwork on the middle strategic neighborhood. Consequently, it is a logically coherent operation to select an equilibrium subnetwork for each strategic neighborhood and then take a “Cartesian product” to obtain the overall network $\bm{A}$ (see §SA.4.2 of Leung, 2020, for a detailed elaboration of this idea).
@@ 4.5 Main Results
$\bar{\Phi}_{\zeta}(\bar{V}^{-1}(\delta,0))=\bar{\Phi}_{\zeta}(\theta_{3}\delta^{2}-\theta_{1}-\theta_{2}),$
@@ 6 Simulation Study
We consider equality tests involving the following the network moment. Define the node statistic $\psi_{i}(\mathcal{N}_{n})=\psi_{i}^{c}-\psi_{i}^{f}$ where
@@
where $c\in(0,1]$ and $\mu$ is the true value of ${\bf E}[\psi_{1}]$, computed using 40k simulation draws. To simulate size, we set $c=1$, and to simulate power, we set $c\in\{0.8,0.9\}$. We consider the following tests at the 5-percent level. To assess the quality of the normal approximation, we use the “oracle” $t$-test with test statistic $(n^{-1}\sum_{i=1}^{n}\psi_{i}-\mu)/\sigma$, where $\sigma$ is the true standard error, computed by taking the standard deviation of the network moment across 40k simulation draws. We then consider the dependence-robust test (20) and randomization test from § 5.1. The randomization test will use data from either 5 or 8 independent networks of identical size, while the other tests only utilize data from one of the networks. We simulate size and power using 5k simulation draws.
@@
The last six columns of Table 1 compare power. The differences between the oracle and dependence-robust tests are due to the latter’s slower rate of convergence. A comparison of the power of the randomization and dependence-robust tests is not apples-to-apples because the former utilizes 5 to 8 times the sample size. However, we see that the randomization test with only 5 networks is extremely conservative and poor in terms of power, performing significantly worse than the dependence-robust test despite having 5 times the sample size. These results confirm the discussion in § 5.
@@ Appendix A Bounded Subnetwork Moments
The proof largely follows that of Theorem 2.1 of Penrose and Yukich (2008) (henceforth PY). We next present several auxiliary lemmas and state the proof at the end of this subsection.
@@
$\begin{aligned} =\frac{e^{-n}n^{m}}{m!}\frac{m!}{m_{1}!m_{2}!}{\bf P}(\bm{X}_{m_{1}}\in S_{1},\,&\bm{T}_{m_{1}}\in\mathcal{U}_{m_{1}},\bm{\zeta}_{m_{1}}\in\mathcal{Z}_{m_{1}},\bm{X}_{m_{1}+1,\dots,n}\in S_{2},\\ &\bm{T}_{m_{1}+1,\dots,n}\in\mathcal{U}_{m_{2}},\bm{\zeta}_{m_{1}+1,\dots,n}\in\mathcal{Z}_{m_{2}})\end{aligned}$
@@
The third equality is true because types are identically distributed, and $N(S_{1})=m_{1}$ and $N(S_{2})=m_{2}$ imply $N_{n}=m$ given that $S_{1},S_{2}$ partition the space. The last line is true because disjoint sets of nodes have independent types and random-utility shocks.
@@
The lemmas that follow utilize the following coupling whose essential idea follows Penrose and Yukich (2003), p. 298. As in § 2.2, define i.i.d. types $\{(X_{i},Z_{i})\}_{i\in\mathbb{N}}$ and i.i.d. random-utility shocks $\{\zeta_{ij}\colon i\neq j,\{i,j\}\subseteq\mathbb{N}\}$, with the two sets mutually independent. For any $n\in\mathbb{N}$, let $N_{n}^{\prime}$ be an independent copy of $N_{n}$ (defined in § 3.2), independent of all other primitives.
@@
The vector $V_{1}(x,y)$ only involves binomial models, whereas $V_{2}(x,y)$ only involves Poissonized models. By the coupling construction, $\mathcal{N}_{x}$ and $\mathcal{N}_{y}$ only involve nodes with labels 5 or greater, and by definition, $\mathcal{N}_{5;\ell_{n}}$ only involves nodes with labels 5 through $\ell_{n}+5$. Then $\psi_{1,x}(\mathcal{N}_{5;\ell_{n}}\cup\{1,2,3\})$, for example, involves nodes 1, 2, 3, and 5 through $\ell_{n}+5$, so it depends on $(X_{1},Z_{1})\equiv(x,z)$ and $(X_{2},Z_{2})\equiv(y,z_{2})$, and $(X_{3},Z_{3})\equiv(x+r_{n}z,z_{3})$, as well as the types of nodes 5 through $\ell_{n}+5$, which are random.
@@
$\lvert{\bf E}^{\star}\left[\big{(}\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1,3\})-\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1\})\big{)}\psi_{2,y}(\mathcal{N}_{5;m-2}\cup\{1,2,3\})\right]\rvert\\ \leq\big{(}\lVert\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1,3\})\rVert_{p}^{\star}+\lVert\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1\})\rVert_{p}^{\star}\big{)}\cdot\lVert\psi_{2,y}(\mathcal{N}_{5;m-2},\cup\{1,2,3\})\rVert_{p}^{\star}\\ \times{\bf P}^{\star}\big{(}\bm{R}(1,\tau_{x,r_{n}}\bm{T}_{\mathcal{N}_{5;\ell-1}\cup\{1,3\}},\bm{\zeta}_{\mathcal{N}_{5;\ell-1}\cup\{1,3\}},\lambda_{\ell+1})>\lVert z\rVert\big{)}^{1-2/p}\leq C\tau_{b,\epsilon}(\lVert z\rVert)^{1-2/p},$
@@
$\bar{p}_{r}(x;x^{\prime})={\bf P}\big{(}\bar{V}(r^{-1}\lVert x-x^{\prime}\rVert,\zeta_{ij})>0\}\big{)}.$ (SA.2.4)
@@
[…]
@@
where we abbreviate $\tilde{\bf E}[\cdot]={\bf E}[\cdot\mid X_{j_{\ell}+2}=x_{j_{\ell}+2},\dots,X_{j_{\ell+\ell}}=x_{j_{\ell+\ell^{\prime}}}]$. By repeatedly applying the Cauchy-Schwarz inequality and Assumption 7(a), we can bound this by
@@
by a change of variables to polar coordinates. Since $\bar{\Phi}_{\zeta}(\bar{V}^{-1}(u,0))\leq c_{3}e^{-c_{4}u}$ for some $c_{3},c_{4}>0$ by Assumption 1, the second term involving $\bar{\Phi}_{\zeta}$ is at least as large as the first term for sufficiently $w$ large, as can be seen from integration by parts. Then
@@
$\cup\big{\{}\max_{i\in Q_{2}(\{1,3\})}r_{n}^{-1}\lVert X_{1}(\bm{t}_{k})-X_{i}(\bm{t}_{k})\rVert>w\big{\}}$ $\cup\big{\{}\max_{i\in Q_{3}(\{1,2\})}r_{n}^{-1}\lVert X_{1}(\bm{t}_{k})-X_{i}(\bm{t}_{k})\rVert>w\big{\}}$
$\cup\big{\{}\max_{i\in Q_{2}(\{1,3\})}r_{n}^{-1}\lVert X_{1}(\bm{t}_{k})-X_{i}(\bm{t}_{k})\rVert>w\big{\}}.$

### v8

@@ 1 Introduction
The next section presents a model of strategic network formation and defines network moments. In § 3, we state high-level conditions for a CLT and outline its proof. Readers interested in the low-level CLT for strategic network formation may skip § 3 for § 4, which presents primitive sufficient conditions. We also outline a general methodology for verifying the high-level condition that can be applied to other network models. In § 5, we discuss practical inference procedures. We present results from a simulation study in § 6, and § 7 concludes. Proofs can be found in the supplemental appendix.
@@ 2.1 Homophily and Sparsity
For any $\delta\in\mathbb{R}_{+}$, $\bar{V}(\delta,\cdot)$ is invertible and increasing in its second argument, and its inverse $\bar{V}^{-1}(\delta,\cdot)$ satisfies $\limsup_{\delta\rightarrow\infty}\delta^{-1}\log\bar{\Phi}_{\zeta}(\bar{V}^{-1}(\delta,0))<0$. Furthermore, there exists $\kappa>0$ such that, for any $n\in\mathbb{N}$,
@@ 2.2 Strategic Interactions and Equilibrium Selection
where $\{S_{n}\}_{n\in\mathbb{N}}$ is a sequence of permutation-invariant functions. [^1] For any bijection $\pi\colon\mathcal{N}_{n}\rightarrow\mathcal{N}_{n}$, we abuse notation by defining $\pi(r_{n}^{-1}\bm{T}_{n})=((r_{n}^{-1}X_{\pi(i)},Z_{\pi(i)}))_{i\in\mathcal{N}_{n}}$ and $\pi(\bm{A})=(A_{\pi(i)\pi(j)})_{i,j}$. We say $S_{n}(\cdot)$ is permutation-invariant if $S_{n}(i,j,r_{n}^{-1}\bm{T}_{n},\bm{A})=S_{n}(\pi(i),\pi(j),\pi(r_{n}^{-1}\bm{T}_{n}),\pi(\bm{A}))$. This means that the function does not directly depend on node labels which have no economic content in our model.
@@ 3.2 Outline of Proof
We prove (13) by adapting a theorem due to Penrose and Yukich (2008) (see our Theorem SA.1.1). We proceed by first partitioning the support of $X_{1}$ into cubes $Q_{1},\dots,Q_{V_{n}}$ of slowly growing volume and representing the moment $n^{-1}\sum_{i=1}^{N_{n}}\psi_{i}(\mathcal{N}_{N_{n}})$ as a double sum over cubes and nodes within cubes $n^{-1}\sum_{i=1}^{V_{n}}\sum_{j=1}^{N_{i}}\psi_{ij}(\mathcal{N}_{N_{n}})$, where $N_{i}$ is the number of nodes positioned in $Q_{i}$ and $ij$ indexes the $j$th node in $Q_{i}$. Spatial independence of the Poisson process implies independence of node positions across cubes. However, node statistics are complex functionals of the process and hence are not generally independent across cubes.
Since stabilization implies that node statistics $\psi_{ij}(\mathcal{N}_{n})$ are primarily determined by nodes relatively proximate to $ij$, and hence cubes relatively proximate to $Q_{i}$, it can be shown that the dependence structure of $\{\sum_{j=1}^{N_{i}}\psi_{ij}(\mathcal{N}_{N_{n}})\}_{i=1}^{V_{n}}$ is “approximately” characterized by a certain “dependency graph.” This is a network in which, roughly speaking, two observations are linked if and only if they are dependent. In our case, observations are cubes, which we connect if and only if they are relatively proximate. Careful construction of the cubes ensures a small approximation error for the dependency graph characterization, and applying a CLT for dependency graphs (Chen and Shao, 2004) delivers the result. For precise details, see § SA.1.1.
@@ 4.1 Strategic Neighborhood
$D_{ij}=\bm{1}\big\{-\theta_{2}<\theta_{1}-\theta_{3}(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert)^{2}+\zeta_{ij}\leq 0\big\}.$
@@ 4.2.1 Best-Response Chains
The previous example pertains to interventions that perturb non-robust potential links, specifically $A_{56}$. Interventions that perturb robust links, such as $A_{67}$, trigger best-response dynamics in all strategic neighborhoods containing the link, which would be the two right-most strategic neighborhoods in Figure 1.
@@ 4.3 Strength of Interactions
$h^{*}(x,z)$ $=\kappa\int_{\mathbb{R}^{2}}\left[\Phi(\theta_{1}+\theta_{2}-\theta_{3}\lVert u\rVert^{2})-\Phi(\theta_{1}-\theta_{3}\lVert u\rVert^{2})\right]\,\text{d}u$
@@
$h^{*}(x,z)=\frac{\kappa\pi}{\theta_{3}}\big(q(\theta_{1}+\theta_{2})-q(\theta_{1})\big).$
@@ 4.4 Coordination
However, if $H$ is specifically a strategic neighborhood, then under Assumption 2, its subnetwork is in fact pairwise stable regardless of the state of the ambient network. Consider Figure 1 and a counterfactual intervention that removes all nodes from the network outside of $C_{5}^{+}=\{2,\ldots,7\}$. The intervention has no impact on the pairwise stability of $A_{23}$ and $A_{67}$ since these are robust links. It also has no impact on the pairwise stability of non-robust potential links between nodes in $C_{5}^{+}$, for example $A_{45}$ since neither nodes 4 nor 5 are linked to $\mathcal{N}_{n}\backslash C_{5}^{+}$ in $\bm{A}$ and strategic interactions are local by Assumption 2. That is, the non-robust potential links in $C_{5}^{+}$ are separated from $\mathcal{N}_{n}\backslash C_{5}^{+}$ by a “buffer” of robust links, so the intervention does not trigger a chain of best-responses that affects $\bm{A}_{C_{5}^{+}}$. This illustrates the following result.
@@
Recall that components of $\bm{D}$ do partition $\mathcal{N}_{n}$, while strategic neighborhoods are obtained by adding nodes that are robustly linked to components. Then the link between any node pair that lies in multiple strategic neighborhoods must necessarily be robust and therefore have the same realization under any pairwise stable equilibrium. That is, any pairwise stable subnetwork on the left strategic neighborhood in Figure 1 sets $A_{23}=1$, as does any pairwise stable subnetwork on the middle strategic neighborhood. Consequently, it is a logically coherent operation to select an equilibrium subnetwork for each strategic neighborhood and then take a “union” to obtain the overall network $\bm{A}$ (see §SA.4.2 of Leung, 2020, for a detailed elaboration of this idea).
@@ 4.5 Main Results
$\bar{\Phi}_{\zeta}(\bar{V}^{-1}(\delta,0))=\bar{\Phi}_{\zeta}(\theta_{3}\delta^{2}-\theta_{1}-\max\{\theta_{2},0\}),$
@@ 6 Simulation Study
We consider equality tests involving the following network moment. Define the node statistic $\psi_{i}(\mathcal{N}_{n})=\psi_{i}^{c}-\psi_{i}^{f}$ where
@@
where $c\in(0,1]$ and $\mu$ is the true value of ${\bf E}[\psi_{1}]$, computed using 40k simulation draws. To simulate size, we set $c=1$, and to simulate power, we set $c\in\{0.8,0.9\}$. We consider the following tests at the 5-percent level. To assess the quality of the normal approximation, we use the “oracle” $t$-test with test statistic $(n^{-1}\sum_{i=1}^{n}\psi_{i}-c\mu)/\sigma$, where $\sigma$ is the true standard error, computed by taking the standard deviation of the network moment across 40k simulation draws. We then consider the dependence-robust test (20) and randomization test from § 5.1. The randomization test will use data from 5, 6, or 8 independent networks of identical size, while the other tests only utilize data from one of the networks. We simulate size and power using 5k simulation draws.
@@
The last six columns of Table 1 compare power. The differences between the oracle and dependence-robust tests are due to the latter’s slower rate of convergence. A comparison of the power of the randomization and dependence-robust tests is not apples-to-apples because the former utilizes 5 to 8 times the sample size. However, we see that the randomization test requires at least 6 networks to obtain nontrivial power. These results confirm the discussion in § 5.
@@ Appendix A Bounded Subnetwork Moments
The proof largely follows that of Theorem 2.1 of Penrose and Yukich (2008) (henceforth PY). We next present several auxiliary lemmas and provide the proof at the end of this subsection.
@@
$\begin{aligned} =\frac{e^{-n}n^{m}}{m!}\frac{m!}{m_{1}!m_{2}!}{\bf P}(\bm{X}_{m_{1}}\in S_{1},\,&\bm{T}_{m_{1}}\in\mathcal{U}_{m_{1}},\bm{\zeta}_{m_{1}}\in\mathcal{Z}_{m_{1}},\bm{X}_{m_{1}+1,\dots,m}\in S_{2},\\ &\bm{T}_{m_{1}+1,\dots,m}\in\mathcal{U}_{m_{2}},\bm{\zeta}_{m_{1}+1,\dots,m}\in\mathcal{Z}_{m_{2}})\end{aligned}$
@@
The third equality holds because types are identically distributed, and $N(S_{1})=m_{1}$ and $N(S_{2})=m_{2}$ imply $N_{n}=m$ given that $S_{1},S_{2}$ partition the space. The last line is true because disjoint sets of nodes have independent types and random-utility shocks.
@@
The lemmas that follow utilize the following coupling whose essential idea is from Penrose and Yukich (2003), p. 298. As in § 2.2, define i.i.d. types $\{(X_{i},Z_{i})\}_{i\in\mathbb{N}}$ and i.i.d. random-utility shocks $\{\zeta_{ij}\colon i\neq j,\{i,j\}\subseteq\mathbb{N}\}$, with the two sets mutually independent. For any $n\in\mathbb{N}$, let $N_{n}^{\prime}$ be an independent copy of $N_{n}$ (defined in § 3.2), independent of all other primitives.
@@
The vector $V_{1}(x,y)$ involves only binomial models, whereas $V_{2}(x,y)$ involves only Poissonized models. By the coupling construction, $\mathcal{N}_{x}$ and $\mathcal{N}_{y}$ involve only nodes with labels 5 or greater, and by definition, $\mathcal{N}_{5;\ell_{n}}$ involves only nodes with labels 5 through $\ell_{n}+5$. Then $\psi_{1,x}(\mathcal{N}_{5;\ell_{n}}\cup\{1,2,3\})$, for example, involves nodes 1, 2, 3, and 5 through $\ell_{n}+5$, so it depends on $(X_{1},Z_{1})\equiv(x,z)$ and $(X_{2},Z_{2})\equiv(y,z_{2})$, and $(X_{3},Z_{3})\equiv(x+r_{n}z,z_{3})$, as well as the types of nodes 5 through $\ell_{n}+5$, which are random.
@@
$\lvert{\bf E}^{\star}\left[\big(\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1,3\})-\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1\})\big)\psi_{2,y}(\mathcal{N}_{5;m-2}\cup\{1,2,3\})\right]\rvert\\ \leq\big(\lVert\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1,3\})\rVert_{p}^{\star}+\lVert\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1\})\rVert_{p}^{\star}\big)\cdot\lVert\psi_{2,y}(\mathcal{N}_{5;m-2}\cup\{1,2,3\})\rVert_{p}^{\star}\\ \times{\bf P}^{\star}\big(\bm{R}(1,\tau_{x,r_{n}}\bm{T}_{\mathcal{N}_{5;\ell-1}\cup\{1,3\}},\bm{\zeta}_{\mathcal{N}_{5;\ell-1}\cup\{1,3\}},\lambda_{\ell+1})>\lVert z\rVert\big)^{1-2/p}\leq C\tau_{b,\epsilon}(\lVert z\rVert)^{1-2/p},$
@@
$\bar{p}_{r}(x;x^{\prime})={\bf P}\big(\bar{V}(r^{-1}\lVert x-x^{\prime}\rVert,\zeta_{ij})>0\big).$ (SA.2.4)
@@
(SA.2.18) $\leq(\ell+\ell^{\prime}+1)\sup_{x_{j_{0}}}\underbrace{\int_{\mathbb{R}^{d}}\cdots\int_{\mathbb{R}^{d}}}_{\ell\text{ times}}\prod_{k=1}^{\ell}\bar{p}_{r_{n}}(x_{j_{k-1}},x_{j_{k}})\,nf(x_{j_{k}})\,\text{d}x_{j_{k}}$
@@
where we abbreviate $\tilde{\bf E}[\cdot]={\bf E}[\cdot\mid X_{j_{\ell}+2}=x_{j_{\ell}+2},\dots,X_{j_{\ell+\ell^{\prime}}}=x_{j_{\ell+\ell^{\prime}}}]$. By repeatedly applying the Cauchy-Schwarz inequality and Assumption 7(a), we can bound this by
@@
by a change of variables to polar coordinates. Since $\bar{\Phi}_{\zeta}(\bar{V}^{-1}(u,0))\leq c_{3}e^{-c_{4}u}$ for some $c_{3},c_{4}>0$ by Assumption 1, the second term involving $\bar{\Phi}_{\zeta}$ is at least as large as the first term for sufficiently large $w$, as can be seen from integration by parts. Then
@@
$\cup\big\{\max_{i\in Q_{2}(\{1,3\})}r_{n}^{-1}\lVert X_{1}(\bm{t}_{k})-X_{i}(\bm{t}_{k})\rVert>w\big\}$ $\cup\big\{\max_{i\in Q_{3}(\{1,2\})}r_{n}^{-1}\lVert X_{1}(\bm{t}_{k})-X_{i}(\bm{t}_{k})\rVert>w\big\}.$